OpenStudy (anonymous):

Find the value of x. Round to the nearest tenth. Answer A. 10.3 B. 31.4 C. 10.7 D. 31.8

OpenStudy (anonymous):



OpenStudy (anonymous):

In respect to angle 35 where is x and side 18? Opposite, adjacent or hypotenuse?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What have you learned so far that will help us solve this problem?

OpenStudy (anonymous):

sine cos tan

OpenStudy (anonymous):

sin=SOH cos=CAH tan=TOA where O=opposite A= Adjacent H=Hypotenuse In respect to angle 35 side 18 and x are opposite and the hypotenuse O and H ok? Tell me which one uses both O and H?

OpenStudy (anonymous):

sin

OpenStudy (anonymous):

Yes \[\sin(35^o)=\frac{opposite}{hypotenuse}\]\[\sin(35^o)=\frac{18}{x}\]\[{x*}\sin(35^o)=18\] \[x=\frac{18}{\sin (35^o)}\]

OpenStudy (anonymous):

10.3

OpenStudy (anonymous):

no 31.3

OpenStudy (anonymous):

Yes you are correct

OpenStudy (anonymous):

Remember sin=SOH cos=CAH tan=TOA

OpenStudy (anonymous):

O=opposite side A=adjacent side H=hypotenuse